# Unveiling In-Context Learning: A Coordinate System to Understand Its Working Mechanism

### v1

@@ Abstract
Large language models (LLMs) exhibit remarkable in-context learning (ICL) capabilities. However, the underlying working mechanism of ICL remains poorly understood. Recent research presents two conflicting views on ICL: One attributes it to LLMs’ inherent ability of task recognition, deeming label correctness and shot numbers of demonstrations as not crucial; the other emphasizes the impact of similar examples in the demonstrations, stressing the need for label correctness and more shots. In this work, we provide a Two-Dimensional Coordinate System that unifies both views into a systematic framework. The framework explains the behavior of ICL through two orthogonal variables: whether LLMs can recognize the task and whether similar examples are presented in the demonstrations. We propose the peak inverse rank metric to detect the task recognition ability of LLMs and study LLMs’ reactions to different definitions of similarity. Based on these, we conduct extensive experiments to elucidate how ICL functions across each quadrant on multiple representative classification tasks. Finally, we extend our analyses to generation tasks, showing that our coordinate system can also be used to interpret ICL for generation tasks effectively.
@@ A Coordinate System to Understand Its Working Mechanism
Jinlan Fu National University of Singapore jinlanjonna@gmail.com                        Xiaoyu Shen Eastern Institute of Technology, Ningbo xyshen@eitech.edu.cn
@@ 1 Introduction
Large language models (LLMs) have demonstrated impressive in-context learning (ICL) capabilities (Brown et al., 2020), i.e., when provided with few-shot examples, LLMs can effectively perform a broad range of tasks without requiring parameter updates (Zhao et al., 2021; Min et al., 2022a; Wei et al., 2023a). The simplicity of this method, combined with its zero training cost and the versatility of applying a single model across various tasks, has made ICL a promising approach to fully leveraging the potential of LLMs.
@@
Figure 1: An overview of the proposed two-dimensional coordinate system for ICL. The y-axis indicates whether the model can recognize the task during ICL, while the x-axis indicates whether similar examples are included in the demonstrations.
@@
However, the underlying working mechanism of ICL remains an open question (Dai et al., 2022; Akyürek et al., 2022). Existing works hold two conflicting views to explain ICL: The first view argues that LLMs implicitly learn tasks required for downstream applications during the pre-training stage, and in-context demonstrations simply provide cues for them to recognize the task Xie et al. (2022). There have been empirical supports for this hypothesis which show that the performance of ICL is insensitive to label correctness of demonstrations, or the number of demonstration shots Min et al. (2022b); Chen et al. (2023); Zhang et al. (2024); Zhu et al. (2024). The second view, on the contrary, suggests that LLMs explicitly learn from similar examples in the demonstrations during the inference stage Liu et al. (2021). Selecting demonstrations similar to the test sample and ensuring their label correctness, or increasing the number of demonstration shots can both improve the downstream performance Rubin et al. (2022); Ye et al. (2023); Levy et al. (2023); Bertsch et al. (2024).
@@
Although both of the above views hold in specific cases, neither can fully explain the working mechanism of ICL from a holistic perspective. In this work, we seek a unified framework encompassing both views to systematically unveil ICL. To do so, we first introduce the peak inverse rank metric to accurately identify the task recognition capability of LLMs. Based on this metric, we observe that LLMs do not always recognize tasks during ICL, even when provided with correct labels in demonstrations. Additionally, the presence of similar examples and the ability of LLMs to recognize tasks are independent variables; fulfilling one does not imply fulfilling the other. Hence, we suggest that the effectiveness of ICL should be described through the interactions of these two orthogonal variables, resulting in four distinct ICL scenarios. To intuitively conceptualize it, we consider each variable as a coordinate axis and visualize all ICL scenarios within a Two-Dimensional Coordinate System, as illustrated in Figure 1. In each quadrant of the coordinate system, we systematically analyze 8 models spanning up to 40B parameters across three major LLM families. We examine a wide range of mainstream classification tasks and have the following main findings:
@@
In general, our findings indicate that the label correctness of similar examples consistently and significantly influences the performance of ICL. When considering scenarios above the x-axis, the label correctness of dissimilar examples can also impact the model’s confidence in recognizing the task. Conversely, when considering scenarios below the x-axis, increasing the number of demonstration shots can markedly enhance ICL performance.
@@ 2.1 The Coordinate Axes
A recent work by Wang et al. (2023) reveals that label words act as semantic anchors, accumulating information of corresponding demonstrations in the shallow layers. The information associated with these anchors is then aggregated in the deeper layers to form the model’s final predictions. Inspired by this, we find that examining whether the hidden states of label tokens at internal layers possess task semantics can serve as an indicator of if ICL has recognized the task. We use a technique called the logit lens (nostalgebraist, 2020; Geva et al., 2021; Dar et al., 2023), which projects transformer representations into the vocabulary space, thereby enabling us to interpret abstruse representations in a human-interpretable manner. Specifically, we project the hidden states of each layer corresponding to the label tokens into the vocabulary space by multiplying them with the pre-trained language modeling head $E$, thus decoding the hidden states of each layer. After obtaining the vocabulary distribution of the hidden states for each layer, we calculate the inverse of the rank of the task-representative token within the vocabulary distribution for each layer. We use the peak inverse rank (PIR) across all layers as our metric to determine whether ICL has recognized the task. A high PIR indicates that ICL has successfully recognized the task, while a low PIR suggests a lack of task understanding capabilities.
@@
Figure 2: The PIR of "capital" at the last label token using Llama-2-7B, before and after replacing all labels with semantically unrelated words.
@@ 3.1 Experimental Settings
We employ neutral delimiters to ensure that the models do not receive any task-specific information from the delimiters. For the specific delimiters used for each dataset, please refer to Appendix E.
@@
We consider an individual prediction correct only when the token with the highest logit within the entire vocabulary matches the ground truth label of the test sample. We measure accuracy by the proportion of correct predictions on the whole dataset.
@@
Figure 6: The average accuracy of zero-shot and ICL with varying numbers of input-label pairs across different models in the second quadrant.
@@ 3.4 Third Quadrant
[…]
Figure 9: The average accuracy of (0,1,6,12)-shot ICL w/o instructions and 1-shot ICL with instructions.
@@ 6 Extension to Generation Tasks
For this reason, we consider an in-context translation task, with all examples in the demonstrations adhering to a strict subject-verb-object structure, as illustrated in Figure 11. If the ICL output also conforms to this strict subject-verb-object structure, it indicates that the tokens "He," "reads," and "book" were generated by performing the explicit sub-tasks of "subject," "verb," and "object," respectively. We analyze the label sentence of the last example in the demonstration. Figure 11 illustrates that for the subject token "He," the label token is "They," with the other tokens rarely generating "subject" semantics. For the verb token "reads," the label token is "watch," with the other tokens seldom generating "verb" semantics. Similarly, for the object token "book," the label token is "TV," and the other tokens rarely generate "object" semantics. This phenomenon of label token sliding within the label sentence provides evidence supporting the plausibility of our hypothesis $H$.
@@ 7 Related Work
Current studies exploring the working mechanisms of ICL can be primarily divided into three streams: theoretical framework, empirical understanding, and logit lens.
@@
Empirical Understanding.
Many studies have explored the influencing factors of ICL, including the order of demonstration examples (Lu et al., 2022), the selection of label words (Min et al., 2022b; Yoo et al., 2022), and the choice of demonstrations (Liu et al., 2021; Rubin et al., 2022). Designing appropriate demonstration construction strategies (Ye et al., 2023; Li et al., 2023) and employing calibration techniques (Zhao et al., 2021; Min et al., 2022a) can significantly enhance ICL performance.
@@
Recently, some studies use the logit lens technique (nostalgebraist, 2020; Geva et al., 2021; Dar et al., 2023) to project complex feature representations into the vocabulary space to study the mechanisms of ICL. Merullo et al. (2023) discover three distinct stages of processing by decoding the next token prediction at each layer. Todd et al. (2024) find that a small number of attention heads transport a compact representation of the demonstrated task. Yu and Ananiadou (2024) provide insights into the mechanisms of ICL in the context of task learning.
@@ Appendix G Impact of Random Label Replacement in the Second Quadrant
We consider two settings: correct input-label mapping and random input-label mapping. In the former, all input-label pair mappings in the demonstration are correct. In the latter, for each input-label pair, the label is randomly selected from the label space. We use $k=6$ in-context examples. The results reflect averages from five random seeds and all datasets in which models can recognize tasks.

### v2

@@ Abstract
Large language models (LLMs) exhibit remarkable in-context learning (ICL) capabilities. However, the underlying working mechanism of ICL remains poorly understood. Recent research presents two conflicting views on ICL: One emphasizes the impact of similar examples in the demonstrations, stressing the need for label correctness and more shots. The other attributes it to LLMs’ inherent ability of task recognition, deeming label correctness and shot numbers of demonstrations as not crucial. In this work, we provide a Two-Dimensional Coordinate System that unifies both views into a systematic framework. The framework explains the behavior of ICL through two orthogonal variables: whether similar examples are presented in the demonstrations (perception) and whether LLMs can recognize the task (cognition). We propose the peak inverse rank metric to detect the task recognition ability of LLMs and study LLMs’ reactions to different definitions of similarity. Based on these, we conduct extensive experiments to elucidate how ICL functions across each quadrant on multiple representative classification tasks. Finally, we extend our analyses to generation tasks, showing that our coordinate system can also be used to interpret ICL for generation tasks effectively.  [^1] Our code is publicly available at: https://github.com/EIT-NLP/2D-Coordinate-System-for-ICL.
@@ A Coordinate System to Understand Its Working Mechanism
Jinlan Fu National University of Singapore jinlanjonna@gmail.com                        Xiaoyu Shen Digital Twin Institute Eastern Institute of Technology, Ningbo xyshen@eitech.edu.cn
@@ 1 Introduction
Large language models (LLMs) have demonstrated impressive in-context learning (ICL) capabilities (Brown et al., 2020), i.e., when provided with few-shot examples, LLMs can effectively perform a broad range of tasks without requiring parameter updates (Zhao et al., 2021; Min et al., 2022a; Su et al., 2022; Wei et al., 2023a). The simplicity of this method, combined with its zero training cost and the versatility of applying a single model across various tasks, has made ICL a promising approach to fully leveraging the potential of LLMs.
@@
Figure 1: An overview of the proposed two-dimensional coordinate system for ICL. The y-axis represents cognition, indicating the model’s ability to recognize tasks during ICL, while the x-axis represents perception, reflecting whether similar examples are included in the demonstrations.
@@
However, the underlying working mechanism of ICL remains an open question (Dai et al., 2022; Akyürek et al., 2022; Olsson et al., 2022; Panwar et al., 2024). Existing works hold two conflicting views to explain ICL: The first view argues LLMs explicitly learn from similar examples in the demonstrations during the inference stage Liu et al. (2021). Selecting demonstrations similar to the test sample and ensuring their label correctness, or increasing the number of demonstration shots can both improve the performance Rubin et al. (2022); Ye et al. (2023); Levy et al. (2023); Bertsch et al. (2024); Liu et al. (2024). The second view, on the contrary, suggests that LLMs implicitly learn tasks required for downstream applications during the pre-training stage, and in-context demonstrations simply provide cues for them to recognize the task Xie et al. (2022). There have been empirical supports for this hypothesis which show that the performance of ICL is insensitive to label correctness of demonstrations, or the number of demonstration shots Min et al. (2022b); Chen et al. (2023); Zhang et al. (2024); Zhu et al. (2024).
@@
Although both of the above views hold in specific cases, neither can fully explain the working mechanism of ICL from a holistic perspective. In this work, we seek a unified framework encompassing both views to systematically unveil ICL. To do so, we first introduce the peak inverse rank metric to accurately identify the task recognition capability of LLMs. Based on this metric, we observe that LLMs do not always recognize tasks during ICL, even when correct labels and similar examples are provided in the demonstrations. Conversely, successful task recognition also does not necessarily require the presence of similar examples and correct labels. Hence, we suggest that the effectiveness of ICL should be described through the interactions of these two orthogonal variables, resulting in four distinct ICL scenarios. To conceptualize this intuitively, we represent each variable along two axes: The x-axis denotes perception, indicating the model’s dependence on similar examples for reference. This mirrors human perception, where recognizing similar observed patterns and drawing analogies is crucial for interpreting new information. The y-axis represents cognition, reflecting the model’s task recognition ability. Similar to human cognition, this involves recognizing the right logic learnt from past tasks and reasoning over it to draw the answer, rather than simply replicating observed patterns. Consequently, all ICL scenarios are visualized within a two-dimensional coordinate system, as depicted in Figure 1. In each quadrant of the coordinate system, we systematically analyze 8 models spanning up to 40B parameters across three major LLM families. We examine a wide range of classification tasks and have the following main findings:
@@
In general, our findings indicate that for similar examples, their label correctness has a consistent and significant impact on ICL, especially in scenarios where the model cannot recognize the task and relies heavily on similar examples for inference (below the x-axis). Increasing the number of demonstration shots substantially improves ICL, as it raises the likelihood of matching similar examples. For dissimilar examples, label correctness primarily affects the model’s confidence in task recognition (above the x-axis), but once the task is properly recognized, the effect becomes marginal.
@@ 2.1 The Coordinate Axes
A recent work by Wang et al. (2023) reveals that label words act as semantic anchors, accumulating information of corresponding demonstrations in the shallow layers. The information associated with these anchors is then aggregated in the deeper layers to form the model’s final predictions. Inspired by this, we find that examining whether the hidden states of label tokens at internal layers possess task semantics can serve as an indicator of if ICL has recognized the task. We use a technique called the logit lens (nostalgebraist, 2020; Geva et al., 2021; Dar et al., 2023), which projects transformer representations into the vocabulary space, thereby enabling us to interpret abstruse representations in a human-interpretable manner. Specifically, we project the hidden states of each layer corresponding to the label tokens into the vocabulary space by multiplying them with the pre-trained language modeling head $E$, thus decoding the hidden states of each layer. After obtaining the vocabulary distribution of the hidden states for each layer, we calculate the inverse of the rank of the task-representative token within the vocabulary distribution for each layer. We use the peak inverse rank (PIR) across all layers as our metric to determine whether ICL has recognized the task. For clarity, we provide the formal definition of PIR in Appendix K. A high PIR indicates that ICL has successfully recognized the task, while a low PIR suggests a lack of task understanding capabilities.
@@
Figure 2: The PIR of "capital" at the last label token using Llama-2-7B, before and after replacing labels.
@@ 3.1 Experimental Settings
We use neutral delimiters to avoid providing task-specific information. See Appendix E for details.
@@
We consider an individual prediction correct only when the token with the highest logit within the entire vocabulary matches the label of the test sample. Accuracy is the proportion of correct predictions across the whole dataset.
@@
Figure 6: Averaged accuracy of zero-shot and ICL with varying numbers of demonstrations in the second quadrant.
@@ 4 Effects of Label Correctness and Demonstration Shot Number
[…]
Figure 9: The average accuracy of (1,6,12)-shot ICL w/o instructions and (0,1)-shot ICL with instructions.
@@ 6 Extension to Generation Tasks
For this reason, we consider an in-context translation task, with all examples in the demonstrations adhering to a strict subject-verb-object structure, as illustrated in Figure 11. If the ICL output also conforms to this strict subject-verb-object structure, it indicates that the tokens "He", "reads", and "book" are generated by performing the explicit sub-tasks of "subject", "verb", and "object", respectively. We analyze the label sentence of the last example in the demonstration. Figure 11 illustrates that for the subject token "He", the label token is "They", with the other tokens rarely generating "subject" semantics. For the verb token "reads", the label token is "watch", with the other tokens seldom generating "verb" semantics. Similarly, for the object token "book", the label token is "TV", and the other tokens rarely generate "object" semantics. This phenomenon of label token sliding within the label sentence provides evidence supporting the plausibility of our hypothesis $H$.
@@ 7 Related Work
Current research on ICL mechanisms mainly falls into the following categories:
@@
Empiricism.
Various factors affecting ICL have been studied, such as the order of examples (Lu et al., 2022), the choice of label words (Min et al., 2022b; Yoo et al., 2022), and the selection of demonstrations (Liu et al., 2021; Rubin et al., 2022). Effective demonstration strategies (Ye et al., 2023; Li et al., 2023) can notably boost ICL performance.
@@
Recently, some studies use the logit lens technique (nostalgebraist, 2020; Geva et al., 2021) to project complex feature representations into the vocabulary space to study the mechanisms of ICL. Merullo et al. (2023) discover three distinct stages of processing by decoding the next token prediction at each layer. Todd et al. (2024) find that a small number of attention heads transport a compact representation of the demonstrated task. Yu and Ananiadou (2024) provide insights into the mechanisms of ICL in the context of task learning.
@@ 8 Conclusion
Acknowledgments
This research / project is supported by the National Research Foundation, Singapore under its Industry Alignment Fund – Pre-positioning (IAF-PP) Funding Initiative. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore. We thank Xingluan (AI Cloud computing service), EIT and IDT High Performance Computing Center for providing computational resources for this project. This work corresponds to Fanghua Ye, Jinlan Fu, and Xiaoyu Shen.
@@ Appendix G Impact of Random Label Replacement in the Second Quadrant
We consider two settings: correct input-label mapping and random input-label mapping. In the former, all input-label pair mappings in the demonstration are correct. In the latter, for each input-label pair, the label is randomly selected from the label space. We use $k=6$ in-context examples without instructions. The results reflect averages from five random seeds and all datasets in which models can recognize tasks.
@@ Appendix I Detailed Instructions for Each Dataset
Appendix K Formal Mathematical Definition of PIR
The PIR metric quantifies a model’s ability to recognize tasks. For a given layer $l$ corresponding to the label token, we begin by projecting the hidden state $h_{l}$ into the vocabulary space by multiplying it with the pre-trained language modeling head $E$. The rank of the task-representative token within this projected distribution is then denoted as $rank_{task}(h_{l},E)$. The PIR is formally defined as:
$PIR=\max\limits_{l}\frac{1}{rank_{task}(h_{l},E)}.$ (1)
Appendix L Additional Experiments on the Reuters-21578 Dataset
To further enhance the comprehensiveness of our study, we conduct additional experiments on the Reuters-21578 dataset (Apt’e et al., 1994), which comprises eight label classes. This dataset is validated through PIR as one in which models are unable to recognize tasks.
Experimental Results.
We employ the same experimental setup as outlined in Section 3.4. The results presented in Table 11 corroborate our existing findings, highlighting the tendency of models to frequently predict the label of the initial example. Furthermore, we confirm the insights discussed in Section 4. Table 12 shows that while 1-shot ICL initially underperforms compared to instruction-based 0-shot ICL, it surpasses it as the shot count increases, and instruction-based 1-shot ICL proves more effective than 1-shot ICL without instructions. These results affirm the efficacy of the two proposed directions for improving ICL performance within the third quadrant.
Table 11: Preference (%) of different models for label tokens at different absolute positions on the Reuters-21578.
Models First Label Second Label Third Label Fourth Label Fifth Label Sixth Label Seventh Label Eighth Label
GPT2-XL 33.1 16.55 15.86 13.79 10.34 10.34 0 0
GPT-J 34.67 13.33 12 6.67 10 23.33 0 0
Llama2-7B 45.65 23.91 10.14 8.7 6.52 5.07 0 0
Llama2-7B-chat 27.4 20.55 10.27 10.96 15.07 15.75 0 0
Llama2-13B 34.75 6.78 14.41 19.49 15.25 9.32 0 0
Llama2-13B-chat 24.66 17.81 14.38 16.44 15.07 11.64 0 0
Table 12: For the Reuters-21578 dataset, the average accuracy (%) of (1, 4, 8, 12)-shot ICL without instructions and (0, 1)-shot ICL with instructions.
Models 1-shot 4-shot 8-shot 12-shot Instruction 0-shot Instruction 1-shot
GPT2-XL 2.67 2.44 4.44 23.78 14 14.44
GPT-J 3.33 5.33 12.67 69.56 18 38.44
Llama2-7B 2 4.67 24.22 50.67 26 39.78
Llama2-7B-chat 3.33 22 42.22 78.67 66 48
Llama2-13B 2 3.78 11.11 83.56 26 27.11
Llama2-13B-chat 4 21.33 37.56 84.67 76 35